# A Transfer Attack to Image Watermarks

### v2

@@ 4.2 Train Surrogate Watermarking Models
Transfer attacks require a set of surrogate watermarking models to produce perturbation for a watermarked image. To improve the transferability of the perturbation produced in our transfer attack, a crucial challenge is to encourage diversity among the surrogate watermarking models. The attacker first collects an image dataset (called surrogate dataset) to train the surrogate watermarking models. The surrogate dataset may have a different distribution from the dataset used to train the target watermarking model. To encourage diversity, we propose to adopt bootstrapping [20], which is a widely used technique to train diverse models in statistics. Specifically, the attacker resamples $m$ subsets from the surrogate dataset, each of which is used to train one of the $m$ surrogate watermarking models. Moreover, the attacker can use different neural network architectures and different watermark lengths for the $m$ surrogate watermarking models.
@@ 4.3 Formulate an Optimization Problem
Select a target watermark for a surrogate decoder:  We use $w_{i}^{t}$ to denote the target watermark for the $i$th surrogate decoder, where $i=1,2,\cdots,m$. We consider the following approaches to select $w_{i}^{t}$.
@@
Inverse-Decode (ID). However, generating random watermarks as target watermarks for surrogate decoders may have an issue. Specifically, for each surrogate decoder, approximately 50% of the bits require to be flipped while optimizing the perturbation, given that each bit of the target watermark is uniformly sampled from $\{0,1\}$. Among these bits to be flipped, some may be very robust to pixel changes and thus large perturbation is required to alter them.
@@
To address this issue, we propose to use the inverse of the watermark decoded by each surrogate decoder as the corresponding target watermark. Specifically, we use $S_{i}$ to denote the $i$th surrogate decoder. For a watermarked image $x_{w}$, $S_{i}$ decodes a watermark from it, denoted as $S_{i}(x_{w})$. We use the inverse of the decoded watermark, i.e., $1-S_{i}(x_{w})$, as the target watermark $w_{i}^{t}$ for the $i$th surrogate decoder $S_{i}$, where $1-S_{i}(x_{w})$ means flipping each bit of $S_{i}(x_{w})$. The intuition of this method is that the watermark decoded by each surrogate decoder from the perturbed image is reversed, and thus it is likely that the watermark decoded by the target decoder from the perturbed image is also reversed, i.e., $T(x_{w}+\delta)$ is close to $1-T(x_{w})$. As a result, the bitwise accuracy between $T(x_{w}+\delta)$ and $w$ is very small since $T(x_{w})$ is close to $w$. We note that such bitwise accuracy may be even smaller than $1-\tau$, which means that the double-tail detector can still detect the perturbed image as AI-generated, watermarked image. We address this challenge by early stopping the process of finding the perturbation $\delta$, which we discuss in Section 4.4.
@@
With the inverse watermark as target watermark, it may result in those bits that are less robust to pixel changes being flipped first during perturbation optimization process. As a result, it can lead to a smaller perturbation compared to random target watermark when the same portion of bits are flipped, as shown in our experiments.
@@
Aggregate $m$ surrogate decoders:  Given a target watermark $w_{i}^{t}$ for the $i$th surrogate decoder, the attacker then finds a perturbation $\delta$ for the watermarked image $x_{w}$ by aggregating the $m$ surrogate decoders. We consider the following two ways to aggregate the $m$ surrogate decoders.
@@
Post-Aggregate (PA). A straightforward way is to apply an existing white-box attack to find a perturbation $\delta_{i}$ based on the $i$th surrogate decoder and target watermark. For instance, given the watermarked image $x_{w}$, $i$th surrogate decoder $S_{i}$, and target watermark $w_{i}^{t}$, the attacker can use the state-of-the-art white-box attack in Jiang et al. [11] to find the perturbation $\delta_{i}$. Formally, the attacker gets $m$ perturbations $\{\delta_{i}\}_{i=1}^{m}$. Then, the attacker can aggregate them as the final perturbation $\delta$, i.e., $\delta=aggr(\delta_{1},\cdots,\delta_{m})$, where $aggr$ is an aggregation method. For instance, $aggr$ can be $Mean$ or $Median$, which takes the mean or median of the $m$ perturbations as the final perturbation $\delta$. However, aggregating the perturbations generated by the surrogate decoders may ruin the patterns in them which are effective to mislead the target decoder before aggregation. As a result, the obtained perturbation achieves limited transferability to the target decoder, as shown in our experiments.
@@
Ensemble-Optimization (EO). To address the challenge of Post-Aggregate, Ensemble-Optimization considers the $m$ surrogate decoders when optimizing the perturbation $\delta$. Specifically, Ensemble-Optimization aims to find a minimum perturbation such that the watermark $S_{i}(x_{w}+\delta)$ decoded by each surrogate decoder for the perturbed image $x_{w}+\delta$ is the same as its corresponding target watermark $w_{i}^{t}$. Formally, we formulate an optimization problem as follows:
$\min_{\delta}\ \|\delta\|_{\infty}$ (1)
$s.t.\ S_{i}(x_{w}+\delta)=w_{i}^{t},\forall i=1,2,\cdots,m$
@@
Ensemble-Optimization finds a shared perturbation $\delta$ such that all the $m$ surrogate decoders output their target watermarks for the perturbed image $x_{w}+\delta$. The generalization ability of the shared perturbation for all surrogate decoders could enhance the transferability to the target decoder, as shown in our experiments. Note that we use $\ell_{\infty}$-norm to measure the magnitude of the perturbation. Our method is also applicable to other metrics such as $\ell_{2}$ or SSIM, as shown in our experiments.
@@ 4.4 Solve the Optimization Problem
$\min_{\delta}\ \|\delta\|_{\infty}$ (2)
$s.t.\ l(S_{i}(x_{w}+\delta),w_{i}^{t})<\epsilon,\forall i=1,2,\cdots,m,$
where $l(\cdot,\cdot)$ denotes a metric to measure the distance between two watermarks. For instance, $l(\cdot,\cdot)$ is the mean square error in our experiments.
The reformulated optimization problem is still challenging to solve due to the high non-linearity of the relaxed constraints. To address this challenge, we further reformulate the optimization problem as follows:
@@
where $r$ is a perturbation budget and $\epsilon$ (called attack strength threshold) is used to ensure that the bitwise accuracy between the watermark decoded by each surrogate decoder for the perturbed image and the corresponding target watermark is high enough to produce an effective perturbation.
We use projected gradient descent (PGD) [13] to solve the optimization problem in Equation 3. To better control the strength of our transfer attack to evade double-tail detector, we consider the second constraint in the optimization problem in Equation 3 as a condition for early stopping. Our detailed algorithm to find the perturbation $\delta$ is shown in Algorithm 1 in Appendix. It first initializes $\delta$ to zero and then compute the gradient of the objective function as defined in Equation 3 with respect to $\delta$. The gradient is then used to update $\delta$ according to gradient descend. If the $\ell_{\infty}$-norm of the updated $\delta$ exceeds $r$, it is then projected onto the $\ell_{\infty}$-norm ball with radius $r$. The algorithm terminates and outputs the final perturbation $\delta$ once the number of optimization iterations or the average bitwise accuracy exceeds a predefined threshold.
@@ 5 Theoretical Analysis
We conduct theoretical analysis about the transferability of our transfer attack. Specifically, given the watermarks decoded by the $m$ surrogate decoders for the perturbed image, we derive both an upper bound and a lower bound for the probability that the watermark decoded by the target decoder for the perturbed image matches with the ground-truth watermark. All our proofs are shown in Appendix.
5.1 Notations

### v4

@@ 4.2 Train Surrogate Watermarking Models
Transfer attacks require surrogate watermarking models to generate perturbations for watermarked images. A key challenge is ensuring diversity among these models to improve transferability. The attacker collects a surrogate dataset, potentially with a different distribution from the target model’s dataset, and uses bootstrapping (Efron & Tibshirani, 1994) to promote diversity. Specifically, $m$ subsets are resampled from the dataset, each used to train one of the $m$ surrogate models. Additionally, the attacker can vary neural network architectures and watermark lengths across the models.
@@ 4.3 Formulate an Optimization Problem
Select a target watermark for a surrogate decoder:  We use $w_{i}^{t}$ to denote the target watermark for the $i$th surrogate decoder, where $i=1,2,\cdots,m$. We consider the following three approaches to select $w_{i}^{t}$.
@@
Inverse-Decode (ID). Generating random watermarks for surrogate decoders poses a challenge, as about 50% bits must be flipped during perturbation optimization, given that each bit of the target watermark is uniformly sampled from $\{0,1\}$. Some bits may be resistant to pixel changes, requiring large perturbations.
@@
To address this, we propose using the inverse of the watermark decoded by each surrogate decoder as the target watermark. Let $S_{i}$ denote the $i$th surrogate decoder, which decodes a watermark from a watermarked image $x_{w}$, denoted as $S_{i}(x_{w})$. We use the inverse, $1-S_{i}(x_{w})$, as the target watermark $w_{i}^{t}$ for $S_{i}$. The intuition is that if the surrogate-decoded watermark is reversed, the target decoder’s output $T(x_{w}+\delta)$ will likely also be close to the inverse of $T(x_{w})$, making the bitwise accuracy between $T(x_{w}+\delta)$ and $w$ very small. We note that such bitwise accuracy may be even smaller than $1-\tau$, which means that the double-tail detector can still detect the perturbed image as AI-generated, watermarked image. We mitigate this issue using early stopping, as discussed in Section 4.4.
@@
With the inverse watermark as target watermark, it may result in those bits that are less robust to pixel changes being flipped first during perturbation optimization process, which can lead to a smaller perturbation compared to random target watermark when the same portion of bits are flipped, as shown in our experiments.
@@
Aggregate $m$ surrogate decoders:  Given a target watermark $w_{i}^{t}$ for the $i$th surrogate decoder, the attacker then finds a perturbation $\delta$ for the watermarked image $x_{w}$ by aggregating the $m$ surrogate decoders. We consider two ways to aggregate the $m$ surrogate decoders.
@@
Post-Aggregate (PA). A simple approach is to apply an existing white-box attack to find a perturbation $\delta_{i}$ for each surrogate decoder. For example, given a watermarked image $x_{w}$, the $i$th surrogate decoder $S_{i}$, and target watermark $w_{i}^{t}$, the attacker can use the white-box attack (Jiang et al., 2023) to obtain $\delta_{i}$. This yields $m$ perturbations $\{\delta_{i}\}_{i=1}^{m}$, which are then aggregated into a final perturbation $\delta$ using a method like mean or median. However, aggregation can disrupt effective patterns in the individual perturbations, limiting transferability to the target decoder, as shown in our experiments.
@@
Ensemble-Optimization (EO). To address the challenge of PA, EO considers the $m$ surrogate decoders when optimizing the perturbation $\delta$. Specifically, EO aims to find a minimum perturbation such that the watermark $S_{i}(x_{w}+\delta)$ decoded by each surrogate decoder for the perturbed image $x_{w}+\delta$ is the same as its corresponding target watermark $w_{i}^{t}$. Formally, we formulate an optimization problem as follows:
$\min_{\delta}\ \|\delta\|_{\infty}\quad s.t.\ S_{i}(x_{w}+\delta)=w_{i}^{t},\forall i=1,2,\cdots,m$ (1)
@@
EO finds a single perturbation $\delta$ that makes all $m$ surrogate decoders produce their respective target watermarks for the perturbed image $x_{w}+\delta$. This unified perturbation improves transferability to the target decoder, as demonstrated in our experiments. We use the $\ell_{\infty}$-norm to measure perturbation size, but our method adapts to other metrics like $\ell_{2}$ or SSIM, as shown in our results.
@@ 4.4 Solve the Optimization Problem
$\min_{\delta}\ \|\delta\|_{\infty}\quad s.t.\ l(S_{i}(x_{w}+\delta),w_{i}^{t})<\epsilon^{\prime},\forall i=1,2,\cdots,m,$ (2)
where $l(\cdot,\cdot)$ denotes a metric to measure the distance between two watermarks. For instance, $l(\cdot,\cdot)$ is the mean square error in our experiments. The reformulated optimization problem is still challenging to solve due to the high non-linearity of the relaxed constraints. To address this challenge, we further reformulate the optimization problem as follows:
@@
where $r$ is a perturbation budget and $\epsilon$ (called sensitivity) is used to ensure that the bitwise accuracy between the watermark decoded by each surrogate decoder for the perturbed image and the corresponding target watermark is high enough to produce an effective perturbation.
We use projected gradient descent (PGD) (Madry et al., 2018) to solve the optimization problem in Equation 3. To control the strength of our transfer attack, we apply the second constraint as an early stopping condition. Algorithm 1 (Appendix) outlines the procedure for finding the perturbation $\delta$. The algorithm initializes $\delta$ to zero, computes gradients of the objective in Equation 3, and updates $\delta$ using gradient descent. If the $\ell_{\infty}$-norm of $\delta$ exceeds $r$, it is projected back onto the $\ell_{\infty}$-norm ball. The algorithm stops when a set iteration limit or bitwise accuracy threshold is reached.
@@ 5 Theoretical Analysis
Given the watermarks decoded by the $m$ surrogate decoders for the perturbed image, we derive both an upper bound and a lower bound for the probability that the watermark decoded by the target decoder for the perturbed image matches with the ground-truth watermark, which quantify the transferability of our transfer attack. Specifically, we derive the following theorem. Other details are shown in Appendix B.
Theorem 1.
For any watermarked image $x_{w}$, perturbation $\delta$ satisfying the constraints in Equation 3, and watermarks decoded by $m$ surrogate decoders for the perturbed image, the probability that the $j$th bit of the watermark decoded by $T$ matches the ground-truth watermark $w$ is bounded as follows:
$Pr(T(x_{w}+\delta)_{j}=w_{j}\mid S_{1}(x_{w}+\delta)_{j},\cdots,S_{m}(x_{w}+\delta)_{j})\geq\begin{cases}\mathop{\max}(\beta_{j}-p_{j},0),&1\leq j\leq n,\\ 0,&n<j\leq n_{t}.\end{cases}$
$Pr(T(x_{w}+\delta)_{j}=w_{j}\mid S_{1}(x_{w}+\delta)_{j},\cdots,S_{m}(x_{w}+\delta)_{j})\leq\begin{cases}1-|p_{j}+\beta_{j}-1|,&1\leq j\leq n,\\ 1,&n<j\leq n_{t}.\end{cases}$
6 Experiments
6.1 Experimental Setup
Datasets:  In our experiments, we utilize three publicly available datasets (Wang et al., 2023; Turc & Nemade, 2022; Images, 2023) generated by Stable Diffusion, Midjourney, and DALL-E 2. The first two datasets are used to train and test the target watermarking models, while the last dataset is used to train the surrogate watermarking model. Each training set contains 10,000 images, and each testing set contains 1,000 images. The details of the datasets are introduced in Appendix H.
Surrogate watermarking models and watermark lengths:  Unless otherwise mentioned, we use HiDDeN (Zhu et al., 2018) and its CNN architecture with a 30-bit watermark length to train surrogate watermarking models. This consistent setup serves two purposes: 1) it helps clearly assess our attack’s transferability when the target and surrogate models have different architectures or watermark lengths, and 2) it allows for better analysis of how the number of surrogate models affects transferability, without the confounding factors of varying architectures or watermark lengths.
Target watermarking models and watermark lengths:  For the target watermarking model, we use different choices of watermarking methods, architectures, and watermark lengths to analyze the transferability of our transfer attack in different scenarios. Specifically, we consider HiDDeN (Zhu et al., 2018), StegaStamp (Tancik et al., 2020), Stable Signature (Fernandez et al., 2023), Smoothed HiDDeN (Jiang et al., 2024), and Smoothed StegaStamp (Jiang et al., 2024) to train target watermarking models. The detailed settings of each watermarking method are shown in Appendix I.
Common post-processing methods:  We consider 4 common post-processing methods. They are JPEG, Gaussian noise, Gaussian blur, and Brightness/Contrast. Details are introduced in Appendix J.
Transfer attacks:  In our experiments, we compare with 5 existing transfer attacks and 1 state-of-the-art purification method for adversarial examples. They are WEvade-B-S (Jiang et al., 2023), AdvEmb-RN18 (An et al., 2024), AdvCls-Real&WM (An et al., 2024), AdvCls-Enc-WM1&WM2, MI-CWA (Chen et al., 2024), and DiffPure (Nie et al., 2022). Details are introduced in Appendix K.
Evaluation metrics:  We use two metrics: evasion rate and average perturbation. Evasion rate is the proportion of perturbed images that evade the target watermark detector. Average perturbation, measured by the $\ell_{\infty}$-norm, is averaged across 1,000 watermarked images in the test set. Following  Jiang et al. (2023), with pixel values normalized from [0, 255] to [-1, 1], we divide the perturbation by 2 to express it as a fraction of the full pixel range.
Parameter settings:  We use adversarial training to train watermarking models since it enhances robustness. The detailed settings of parameters are shown in Appendix L
6.2 Experimental Results
Our transfer attack is successful:  Figure 2 and Figure 3 show the evasion rate and average perturbation of our attack, across different numbers of surrogate models, target model architectures, and watermark lengths. First, our attack successfully evades watermark-based detection, achieving an evasion rate over 85% with an average perturbation below 0.25 when using 100 surrogate models, regardless of target architecture or watermark length. Second, more complex target architectures require more surrogate models and higher perturbation. For example, 40 surrogate models and perturbation of 0.09 achieve 100% evasion for a CNN with 30-bit watermarks, while 60 models and 0.14 perturbation are needed for a ResNet. Third, longer/shorter watermarks also increase the attack’s complexity. For a ResNet with 30-bit watermarks, 40 surrogate models and perturbation of 0.08 yield 85% evasion, while 100 models and 0.15 perturbation are needed for 64-bit watermarks.
Our transfer attack outperforms common post-processing:  Figure 4 compares the average perturbations of common post-processing methods and our transfer attack when they achieve the same evasion rates for different target models. For each evasion rate achieved by our transfer attack, we adjust the parameters of the post-processing methods to match it. We find that our attack introduces much smaller perturbations than post-processing methods for the same evasion rate. Note that Brightness/Contrast only reaches up to 75% or 50% evasion for some models, so its curves are shorter in the graphs. For a broader comparison, Figures 8 and 9 in Appendix show results using $\ell_{2}$-norm and SSIM. Our attack consistently introduces smaller perturbations and maintains better visual quality than common post-processing, regardless of the metric used.
Our transfer attack outperforms existing ones:  Figure LABEL:othertf-evasion compares the evasion rates of existing methods and our transfer attack when the target model is ResNet with varying watermark lengths. Since AdvCls-Real&Wm and AdvCls-Enc-WM1&WM2 achieve average perturbations around 0.1 regardless of the budget $r$, we introduce a variant of our method with $r=0.1$ for comparison. For fairness, we constrain MI-CWA and DiffPure to an average SSIM of 0.9, similar to our attack with 100 surrogate models. Other attacks have $r$ set to 0.25. We also report the evasion rate for each method across a wide range of SSIM constraints in Figure 11 in Appendix.
Our transfer attack significantly outperforms existing methods, which mostly achieve 0% evasion, except AdvCls-Enc-WM1&WM2, which reaches 10% on Stable Diffusion with 20-30 bit watermarks but requires access to the target encoder. MI-CWA and DiffPure show low evasion rates due to the limitations of classifier-based attacks and heavy image modification for watermark removal. In contrast, our transfer attack achieves much higher evasion rates while maintaining high image quality, with an average SSIM above 0.97 for $r=0.1$ and above 0.92 for $r=0.25$.
Figure 5: (a) Comparing evasion rates of existing and our transfer attacks. The target model is ResNet and uses watermarks with different lengths. Dataset is Stable Diffusion. Similar results for Midjourney are shown in Figure 10 in Appendix. (b) Evasion rates and (c) average $\ell_{\infty}$-norm perturbation of our transfer attacks to different target watermarking methods.
Our transfer attack transfers to other target watermarking methods:  Figures LABEL:otherwm-evasion and LABEL:otherwm-linf show the evasion rates and average $\ell_{\infty}$-norm perturbations of our transfer attack across different target watermarking methods. To improve transferability, we use equal numbers of HiDDeN and StegaStamp as surrogate models. Our attack remains effective even when surrogate and target methods differ, achieving over 77% evasion across all five target methods with 100 surrogate models. Despite Smoothed HiDDeN and Smoothed StegaStamp being certifiably robust, they are still vulnerable to our attack. While evasion rates are slightly lower for Smoothed HiDDeN with 20 models and Smoothed StegaStamp with over 60 models, we still exceed 77% evasion with up to 100 models, as our perturbations exceed their certified bounds while preserving image quality.
Impact of sensitivity $\epsilon$:  Figures LABEL:diff-epsilon-evasion and LABEL:diff-epsilon-linf in Appendix show the evasion rate and average $\ell_{\infty}$-norm perturbation for different sensitivity values ($\epsilon$) in our transfer attack on Stable Diffusion, with a CNN target model using 30-bit watermarks. Figures LABEL:diff-epsilon-l2 and LABEL:diff-epsilon-ssim in Appendix show results for $\ell_{2}$-norm and SSIM. We observe that fewer surrogate models are needed for the same evasion rate when $\epsilon$ is smaller. However, with very small $\epsilon$ (e.g., 0.1), the evasion rate first increases but then decreases as more surrogate models are used, due to reduced bitwise accuracy in watermark decoding. Across different $\epsilon$ values, the average perturbation (whether measured by $\ell_{\infty}$-norm, $\ell_{2}$-norm, or SSIM) remains nearly constant for a given evasion rate, reflecting the strong correlation between perturbation size and evasion success.
Different variants of our transfer attack:  Our attack has two steps, each with multiple design options. In the first step, we use RD, RS, or ID; in the second step, PA or EO is used. Variants are denoted by combining these symbols, e.g., RD-PA. When PA is used, “Mean” or “Median” indicates the aggregation rule, such as RD-PA-Mean. Figures LABEL:diff-var-evasion and LABEL:diff-var-linf in Appendix show evasion rates and $\ell_{\infty}$-norm perturbations for different variants on Stable Diffusion, with a CNN using 30-bit watermarks. Figures LABEL:diff-var-l2 and LABEL:diff-var-ssim in Appendix show $\ell_{2}$-norm and SSIM results. EO-based variants successfully evade detection, while PA-based variants do not, showing EO is more effective for surrogate decoder aggregation. Among EO variants, ID-EO outperforms RD-EO and RS-EO, achieving 100% evasion with fewer surrogate models, lower perturbations, and higher SSIM.
Figure 6: Comparing empirical bitwise accuracy and theoretical bounds of our transfer attack when the target model is (a) CNN and (b) ResNet. (c) Comparing evasion rates of non-learning-based watermarking models and HiDDeN under Gaussian noise with varying standard deviations.
Theoretical vs. empirical results:  Figures LABEL:theory-cnn and LABEL:theory-resnet show the empirical bitwise accuracy between the watermark decoded by the target decoder and the ground truth, along with our theoretical upper and lower bounds on the Stable Diffusion dataset. To estimate these bounds, we measured unperturbed similarity, transfer similarity, and attack strength using 1,000 test images. Using Theorem 1, we computed the probability for each bit $p_{j}$, then averaged the bounds across all bits. We observe that the empirical results generally fall within our theoretical bounds, with a few exceptions due to the use of only 1,000 images for estimation, leading to minor inaccuracies.
7 Discussion and Limitations
Non-learning-based watermarking:  We focus on transfer attacks against learning-based watermarking models because they are more robust than non-learning-based ones. For example, common post-processing like Gaussian noise can easily remove watermarks from non-learning-based models, but learning-based models resist such attacks (Jiang et al., 2023). Figure LABEL:non-learning-based presents the evasion rates of Tree-Ring (Wen et al., 2024), RingID (Ci et al., 2025), Gaussian Shading (Yang et al., 2024), and HiDDeN with ResNet under Gaussian noise. For the standard deviation larger than 0.1, the non-learning-based methods exhibit significantly higher evasion rates, whereas the learning-based method maintains an evasion rate close to zero. These results underscore the superior robustness of learning-based watermarking methods compared to the non-learning-based ones.
Our attack can also be applied to non-learning-based methods. Figure 15 in the Appendix shows that our attack achieves high evasion rates when the target watermarking method is DWT-DCT (Al-Haj, 2007) while the surrogate models are trained under our default experimental settings. Unlike learning-based methods, DWT-DCT is a non-learning-based watermarking method used by Stable Diffusion. It employs the Discrete Wavelet Transform (DWT) to decompose an image into frequency sub-bands, applies the Discrete Cosine Transform (DCT) to blocks within selected sub-bands, and embeds the watermark by modifying specific frequency coefficients. The watermarked image is then reconstructed using inverse transforms. The result shows that our attack requires more surrogate models compared to learning-based methods. This is because non-learning-based watermarking method is substantially different from learning-based, and our attack needs more surrogate models to increase diversity to enhance transferability.
We acknowledge that the victim may design a new watermarking method entirely different from those used by the attacker as surrogates, and the effectiveness of our attack in such scenarios is unclear. We believe this can be an interesting future work to explore.
Computational cost:  Our transfer attack trains surrogate watermarking models and optimizes perturbations for watermarked images. Thus, the computational cost consists of training time (to train surrogate models) and inference time (to optimize a perturbation for an image). Table 2 in Appendix compares the computational costs of our attack with six baselines. Attacking high-capacity watermarks (Zhong & Shih, 2019) may require more surrogate models, further increasing training cost—a practical limitation of our approach. However, training multiple surrogate models can be parallelized across GPUs, and this training is a one-time process, which helps mitigate the limitation. We acknowledge that our inference time exceeds most baselines but remains acceptable with adequate computational resources.
8 Conclusion and Future Work
In this work, we find that watermark-based detection of AI-generated images is not robust to transfer attacks in the no-box setting. Given a watermarked image, an attacker can remove the watermark by adding a perturbation to it, where the perturbation can be found by ensembling multiple surrogate watermarking models. Our results show that transfer attack based on surrogate watermarking models outperforms those based on surrogate classifiers that treat a watermark-based detector as a conventional classifier. Moreover, leveraging surrogate watermarking models enables us to perform a rigorous analysis on the transferability of our attack. Interesting future work includes extending our transfer attack to text and audio watermarks, as well as designing more robust watermarks.
9 Acknowledgment
We sincerely thank all the reviewers for their constructive feedback. This work was supported by NSF grants No. 2414406, 2131859, 2125977, 2112562, and 1937787, as well as ARO grant No. W911NF2110182.
Appendix A Ethics Concerns
Our proposed method has the potential to compromise the identity and authenticity of AI-generated images, raising ethical and societal concerns. Such misuse could erode trust in digital media by enabling unauthorized modifications or misattributions. To mitigate these risks, it is crucial to develop robust watermarking methods that can withstand our transfer attack. As discussed in Section 7, our attack’s effectiveness is unclear when the target watermarking method is entirely unknown and significantly different from the surrogate methods used by attackers. Therefore, designing novel watermarking techniques with fundamentally different mechanisms and maintaining their confidentiality could serve as an effective strategy to mitigate these risks.
Appendix B Theoretical Analysis
We derive both an upper bound and a lower bound of the probability because we consider double-tail detectors. Specifically, they enable us to compute the upper bound and lower bound of the expected bitwise accuracy between the watermark decoded by the target decoder for the perturbed image and the ground-truth watermark. If such upper bound and lower bound of the expected bitwise accuracy fall between $\tau$ and $1-\tau$, the perturbed image produced by our transfer attack is expected to evade the target watermark-based detector. All our proofs are shown in Appendix.
B.1 Notations